### v1

Abstract
We present new convergence estimates of generalized empirical interpolation methods in terms of the entropy numbers of the parametrized function class. Our analysis is transparent and leads to sharper convergence rates than the classical analysis via the Kolmogorov $n$-width. In addition, we also derive novel entropy-based convergence estimates of the Chebyshev greedy algorithm for sparse $n$-term nonlinear approximation of a target function. This also improves classical convergence analysis when corresponding entropy numbers decay fast enough.
keywords:
Empirical interpolation method, reduced basis greedy algorithm, Chebyshev greedy algorithm, metric entropy numbers, parametrized PDE, dictionary approximation.
{MSCcodes}
41A46, 41A65, 65J05, 65M12
1 Introduction
Parametrized Partial Differential Equations (PDEs) are important mathematical models arising in a variety of physical contexts, e.g, material design, shape optimization, uncertainty quantification, inverse problems etc. For these real-world applications, it is necessary to efficiently evaluate the query of PDE solutions for many instances of the parameter, such as a family of elasticity and diffusion coefficients in deterministic PDEs and different realizations of random fields in stochastic PDEs. In particular, the numerical solver for a parametrized family of PDEs is expected to be much faster than repeatedly running general-purpose PDE solvers, e.g., finite element and finite difference methods, for each parameter instance.
When solving parametric PDEs, model reduction is a ubiquitous strategy that replaces the large-scale high-fidelity PDE model with an easy-to-solve reduced model. In recent decades, the Reduced Basis Method (RBM) is possibly one of the most notable model reduction techniques, see, e.g., [22, 6, 29] and references therein. However, the efficiency of RBMs and other projection-based model reduction methods hinges on the affine parametrization structure which might not be available in practice, see Section 2.1 for more details.
To overcome this drawback, the Empirical Interpolation Method (EIM) was invented in [1] for constructing affinely parametric approximants to coefficient functions in parametrized PDEs. Since then, the EIM was generalized in many ways and becomes an extremely popular tool for obtaining the affine structure for efficient model reduction, see, e.g., [28, 11, 4, 13, 23, 27]. Convergence analysis of EIMs could be found in [28, 5, 12, 24]. In fact, the EIM can be viewed as a weak reduced basis greedy algorithm with varying threshold constants at each iteration. Therefore, the work [24] modified the error analysis of RBMs in [8] and derived convergence rates of the EIM in terms of the Kolmogorov $n$-width $d_{n}(K)$ of the compact set $K$ of parametrized functions. For example, Corollary 14 in [24] shows that for each $\alpha>0$,
$d_{n}(K)\leq Cn^{-\alpha}\Longrightarrow\sup_{f\in K}\|f-\Pi_{n}f\|\leq C(1+\Lambda_{n})^{3}n^{-\alpha+1},$ (1.1)
where $\Pi_{n}$ is the EIM interpolation at the $n$-th iteration, $C$ is an absolute constant, and $\Lambda_{n}$ is the norm of $\Pi_{n}$ and is called the Lebesgue constant.
In this paper, we shall investigate the convergence of EIMs in a way different from [24]. Instead of the Kolmogorov $n$-width, the key concept in our analysis is the metric entropy numbers $\varepsilon_{n}({\rm co}(K))$ of the symmetric convex hull ${\rm co}(K)$ of the underlying function set $K$. Such an idea originates from the author’s work [18] about entropy-based convergence rates of greedy RBMs. Our new error analysis of EIMs is transparent and sharper than classical results in many cases. In Banach spaces, the proposed convergence estimate is
$\sup_{f\in K}\|f-\Pi_{n-1}f\|\leq C(1+\Lambda_{n-1})\left(\prod_{k=1}^{n-1}(1+\Lambda_{k})\right)^{\frac{1}{n}}n\varepsilon_{n}({\rm co}(K)).$ (1.2)
Here we do not make any assumption on $\{\Lambda_{n}\}_{n\geq 0}$ and $\{\varepsilon_{n}({\rm co}(K))\}_{n\geq 0}$. For a standard parametrized compact set $K$, it is shown in [33] that $n^{\frac{1}{2}}\varepsilon_{n}({\rm co}(K))=d_{n}(K)=O(n^{-t})$ for some $t>0$. It is also numerically observed (see [25]) that $\Lambda_{n}=O(n^{\beta})$ for some $\beta>0$. Therefore, (1.2) is sharper than the classical result (1.1) for a wide range of function classes. In addition, the estimate (1.2) could be improved in Hilbert spaces and special Banach spaces, see Section 3 for the detailed analysis and comparisons.
The EIMs are greedy algorithms designed to simultaneously approximate a compact collection of parameter-dependent functions. On the other hand, greedy algorithms are also widely used for sparse nonlinear approximation of a single target function. These algorithms adaptively select basis functions $\{g_{i}\}_{1\leq i\leq n}$ from a compact dictionary $K\subset X$ and use a $n$-term linear combination
$f_{n}=\sum_{i=1}^{n}c_{i}g_{i},\quad g_{i}\in K$
to approximate a function $f\in X$. Important applications of sparse greedy approximations include nonparametric statistical regression [14, 17], matching pursuit in compressed sensing [26], training shallow neural networks [32] etc. We refer to [37, 35] for a thorough introduction to greedy dictionary approximation. Among such greedy algorithms, it is recently shown in [32, 18] that the Orthogonal Greedy Algorithm (OGA) in Hilbert spaces exhibits the fastest convergence rate provided ${\rm co}(K)$ has small entropy numbers.
The OGA was generalized to Banach spaces and named as the Chebyshev Greedy Algorithm (CGA) by Temlyakov in [36]. The convergence of the CGA in its simplest form (see [36]) reads
$\|f-f_{n}\|=O(n^{-1+\frac{1}{s}}),\quad f\in{\rm co}(K),$ (1.3)
where $s\in(1,2]$ is the power in the modulus of smoothness of the underlying Banach space $X$ (see [20, 10, 36]). The second main contribution of our work is the novel convergence estimate of the CGA based on entropy numbers:
$\|f-f_{n}\|\leq C_{f}n^{\frac{1}{s}-\frac{1}{2}}\delta_{X,n}\varepsilon_{n}({\rm co}(K)),$ (1.4)
where $C_{f}$ is an explicit uniform constant depending on the target function $f$, and $\delta_{X,n}$ is the supremum of the Banach–Mazur distance between $\ell_{2}^{n}$ and $n$-dimensional subspaces of $X$. Our convergence estimate (1.4) is a Banach version of the result about OGAs in [18]. In addition, (1.4) leads to improved convergence rates of the CGA under the condition $\delta_{X,n}\varepsilon_{n}({\rm co}(K))=o(n^{-\frac{1}{2}})$, which is true for popular dictionaries in Banach spaces, e.g., the $\texttt{ReLU}_{m}$ dictionary in $L_{p}(\Omega)$ with $1<p<\infty$, see Section 4 for details.
The rest of this paper is organized as follows. In Section 2, we briefly introduce the EIM and the CGA in Banach spaces. In Section 3, we present a convergence analysis of the EIM in terms of entropy numbers. Section 4 is devoted to an entropy-based convergence estimate of the CGA and related numerical experiments.
2 Greedy Algorithms
Let $K$ be a compact set in a Banach space $X$ under the norm $\|\bullet\|=\|\bullet\|_{X}$. By ${\rm co}(K)$ we denote the symmetric convex hull of $K$, i.e.,
$\text{co}(K):=\overline{\Big{\{}\sum_{i}c_{i}g_{i}:\sum_{i}|c_{i}|\leq 1,~{}g_{i}\in K\text{ for each }i\Big{\}}}.$
The Kolmogorov $n$-width of $K$ is defined as
$d_{n}(K)=d_{n}(K)_{X}:=\inf_{X_{n}}\sup_{f\in K}{\rm dist}(f,X_{n}),$
where the infimum is taken over all $n$-dimensional subspaces of $X$, and
${\rm dist}(f,X_{n}):=\inf_{g\in X_{n}}\|f-g\|.$
Convergence of many classical greedy algorithms are analyzed in terms of $d_{n}(K)$, most notably the reduced basis greedy algorithm (see [3, 8, 40]). As mentioned before, we shall follow an alternative way and analyze the convergence of the EIM and the CGA using the entropy numbers of ${\rm co}(K)$ (see [21]):
$\varepsilon_{n}({\rm co}(K))=\varepsilon_{n}({\rm co}(K))_{X}:=\inf\{\varepsilon>0:{\rm co}(K)\text{ is covered by }2^{n}\text{ balls of radius }\varepsilon\}.$
The sequence $\{\varepsilon_{n}({\rm co}(K))\}_{n\geq 0}$ describes the massiveness of ${\rm co}(K).$ It is well-known that $\lim_{n\to 0}d_{n}(K)=0$ and $\lim_{n\to 0}\varepsilon_{n}({\rm co}(K))=0$ for a compact $K.$
2.1 Reduced Basis Greedy Algorithm
In the analysis of RBMs, $K$ is the solution manifold of a parametrized PDE. Let $\mathcal{P}\subset\mathbb{R}^{d}$ be a compact set of parameters. For simplicity of presentation, we consider the boundary value problem for $\mu\in\mathcal{P}$:
$-\nabla\cdot(a_{\mu}\nabla u_{\mu})+b_{\mu}u_{\mu}$ $=F\quad\text{ in }\Omega,$ (2.1a)
$u_{\mu}$ $=0\quad\text{ on }\partial\Omega,$ (2.1b)
where $\Omega\subset\mathbb{R}^{q}$ is the physical domain, $F\in H^{-1}(\Omega)$, and
$0<\inf_{\mu\in\mathcal{P}}\inf_{\Omega}a_{\mu}\leq\sup_{\mu\in\mathcal{P}}\sup_{\Omega}a_{\mu}<\infty,$
$0<\inf_{\mu\in\mathcal{P}}\inf_{\Omega}b_{\mu}\leq\sup_{\mu\in\mathcal{P}}\sup_{\Omega}b_{\mu}<\infty.$
In this case, the solution manifold
$K=\{u_{\mu}\in H_{0}^{1}(\Omega):\mu\in\mathcal{P}\}$
is also compact, and the corresponding RBM approximately solves (2.1) for many instances of the parameter $\mu$ in a low-dimensional subspace. The abstract greedy RBM is formulated in Algorithm 2.1.
Algorithm 2.1 Weak Reduced Basis Greedy Algorithm
Input: a compact set $K\subset X$, an integer $N$, and a sequence $\{\alpha_{n}\}\subset(0,1]$;
Initialization: set $X_{0}=\{0\}$;
[…]
      Step 1: select $f_{n}\in K$ such that
      Step 2: set $X_{n}={\rm span}\{f_{1},\ldots,f_{n}\}$;
EndFor
Output: the subspace $X_{N}$ that approximates $K$.
In RBMs for parametrized PDEs, the threshold sequence $\{\alpha_{n}\}$ in Algorithm 2.1 is indeed constant, i.e., $\alpha_{1}=\cdots=\alpha_{N}$. The output $X_{N}$ is a low-dimensional subspace over which Galerkin solutions for many instances of the parameter are computed during the online stage of the RBM. We refer to [30, 16, 3, 8, 18, 19] for more details about the implementation and analysis of RBMs.
The efficiency of RBMs relies on the affine structure of coefficients, e.g., when $a_{\mu},b_{\mu}$ in (2.1) are of the special form
$a_{\mu}(x)=a_{0}(x)+\sum_{j=1}^{J}\theta_{j}(\mu)a_{j}(x),$
$b_{\mu}(x)=b_{0}(x)+\sum_{m=1}^{M}\omega_{m}(\mu)b_{m}(x).$
Such affine decomposition of $a_{\mu},b_{\mu}$ is not necessarily available in real-world PDE models. To improve the applicability of RBMs, the works [1, 11, 23] etc. have developed a variety of EIMs for approximating parameter-dependent coefficients, e.g., $\{a_{\mu}\}_{\mu\in\mathcal{P}}$ and $\{b_{\mu}\}_{\mu\in\mathcal{P}}$, by affinely separable functions uniformly with respect to $\mu\in\mathcal{P}$. In this paper, we consider the generalized EIM (see [1, 23, 24]) given in Algorithm 2.2.
Algorithm 2.2 Generalized Empirical Interpolation Method
Input: a compact set $K\subset X$, a collection $\mathcal{L}$ of bounded linear functionals acting on $\text{span}\{K\}$, and an integer $N\geq 1$;
Initialization: let $X_{0}=\{0\}$, $\ell_{0}=0\in\mathcal{L}$ and $\Pi_{0}:X\rightarrow X_{0}$ be the zero map;
[…]
      Step 1: select $f_{n}\in K$ such that
      Step 2: construct $X_{n}={\rm span}\{f_{1},\ldots,f_{n}\}$;
      Step 3: select $\ell_{n}\in\mathcal{L}$ such that
      Step 4: let $g_{n}:=r_{n}/\ell_{n}(r_{n})$ and $B_{n}=(\ell_{i}(g_{j}))_{1\leq i,j\leq n}$;
      Step 5: define the interpolation $\Pi_{n}:\text{span}\{K\}\rightarrow X_{n}$ as
EndFor
Output: the empirical interpolation $\Pi_{N}:\text{span}\{K\}\rightarrow X_{N}$.
Under practical assumptions (see [1]), Algorithm 2.2 is well-defined and satisfies the followings: (i) $B_{n}$ is a lower triangular matrix with unit diagonal entries; (ii) $g_{1},\ldots,g_{n}$ is a basis of $X_{n}$; (iii) $\ell_{1},\ldots,\ell_{n}$ is unisolvent on $X_{n}$, and thus $\Pi_{n}^{2}=\Pi_{n}$. Step 5 ensures that $\Pi_{n}$ is an interpolation based on the following degrees of freedom:
$\ell_{i}(\Pi_{n}f)=\ell_{i}(f),\quad i=1,2,\ldots,n,\quad f\in K.$
In the original work [1], $X=L_{\infty}(\Omega)$ and $\mathcal{L}$ corresponds to function evaluations at a set of sample points $\{x_{i}\}_{1\leq i\leq L}$ in the domain $\Omega$, i.e., $\ell_{x_{i}}(f)=f(x_{i})$ for each $\ell_{x_{i}}\in\mathcal{L}$. In practice, $K$ is a collection of functions smoothly parametrized by a compact set of parameters. Introducing basis functions $(h_{1},\ldots,h_{n}):=(g_{1},\ldots,g_{n})B_{n}^{-1},$ the EIM interpolant has the explicit form
$\Pi_{n}f=\ell_{1}(f)h_{1}+\cdots+\ell_{n}(f)h_{n}.$ (2.2)
For the problem (2.1), the compact set $K$ in Algorithm 2.2 is
$K=\{a_{\mu}\in L_{\infty}(\Omega):\mu\in\mathcal{P}\}\text{ or }K=\{b_{\mu}\in L_{\infty}(\Omega):\mu\in\mathcal{P}\}.$
In this case, the coefficients $c_{1},\ldots,c_{n}$ in Step 5 of Algorithm 2.2 depends on $\mu$. The EIM interpolants of $a_{\mu}$ and $b_{\mu}$ in (2.1) are of the form
$(\Pi_{N}a_{\mu})(x)=\sum_{i=1}^{N}c_{i}(\mu)a_{\mu_{i}}(x),$
$(\Pi_{N}b_{\mu})(x)=\sum_{i=1}^{N}c_{i}(\mu)b_{\mu_{i}}(x),$
where $\mu\in\mathcal{P}$ and $x\in\Omega$ are separated. As a result, a RBM for the approximate model
$-\nabla\cdot\big{(}(\Pi_{N}a_{\mu})\nabla\tilde{u}_{\mu}\big{)}+(\Pi_{N}b_{\mu})\tilde{u}_{\mu}$ $=F\quad\text{ in }\Omega,$
$\tilde{u}_{\mu}$ $=0\quad\text{ on }\partial\Omega$
is able to achieve much higher efficiency than a RBM directly applied to (2.1).
2.2 Chebyshev Greedy Algorithm
In the end of this section, we consider another class of greedy algorithms for dictionary approximation. Famous examples of these greedy algorithms include the projection pursuit in statistical regression and the matching pursuit in compressed sensing. Among those algorithms, the OGAs (see Algorithm 2.3) in Hilbert spaces (also known as the orthogonal matching pursuit in compressed sensing) are perhaps the ones with fastest convergence rate.
Algorithm 2.3 Orthogonal Greedy Algorithm
Input: a Hilbert space $H$, a compact set $K\subset H$, a target function $f\in H,$ and an integer $N$;
Initialization: set $f_{0}=0\in H;$
[…]
      Step 1: compute the optimizer
      Step 2: set $H_{n}={\rm span}\{g_{1},\ldots,g_{n}\}$ and compute
      where $P_{H_{n}}$ is the orthogonal projection onto $H_{n}$;
EndFor
Output: the iterate $f_{N}$ that approximates $f$.
Convergence of the OGA have been analyzed in, e.g., [2, 32, 18]. In a Banach space, the CGA is a natural generalization of the OGA, see [36, 7]. Compared with the OGA, greedy approximation in non-Hilbert spaces like $L_{p}$ with $p<2$ is more robust with respect to noise, see [15, 10]. In [38], the CGA is also a theoretical tool for proving Lebesgue-type inequalities.
For an element $g\in X$, its peak functional $F_{g}$ (see [10, 36]) is a bounded linear functional in $X^{\prime}$ that satisfies
$\|F_{g}\|_{X^{\prime}}=1,\quad F_{g}(g)=\|g\|.$ (2.3)
The existence of $F_{g}$ is ensured by the Hahn–Banach extension theorem. The weak version of the CGA (see [36]) is described in Algorithm 2.4.
Algorithm 2.4 Weak Chebyshev Greedy Algorithm
Input: a compact set $K\subset X$, a target function $f\in X$, an integer $N$, and a sequence $\{\alpha_{n}\}_{n\geq 1}\subset(0,1]$;
Initialization: set $f_{0}=r_{0}=0\in X;$
[…]
      Step 1: construct the peak functional $F_{r_{n-1}}\in X^{\prime}$ for $r_{n-1}$;
      Step 2: select $g\in K$ such that
      Step 3: set $X_{n}={\rm span}\{f_{1},\ldots,f_{n}\}$;
      Step 4: select $f_{n}\in X_{n}$ such that
EndFor
Output: the iterate $f_{N}$ that approximates $f$.
When $\alpha_{1}=\cdots=\alpha_{N}=1$ and $X$ is a Hilbert space, the weak CGA reduces to the OGA in Algorithm 2.3.
3 Convergence of the EIM
In this section, we derive a novel convergence estimate for the generalized EIM (Algorithm 2.2). We start with the norm or the Lebesgue constant of the interpolation $\Pi_{n}$ defined by
$\Lambda_{n}=\|\Pi_{n}\|_{K^{\prime}}:=\sup_{0\neq g\in\text{span}\{K\}}\|\Pi_{n}g\|/\|g\|.$
For any $f\in K$, it holds that
$\inf_{g\in X_{n}}\|f-g\|\leq\|f-\Pi_{n}f\|.$ (3.1)
On the contrary, for any $g\in X_{n}$ we have $\|f-\Pi_{n}f\|=\|(I-\Pi_{n})(f-g)\|$ and obtain
$\|f-\Pi_{n}f\|\leq\gamma_{n}\inf_{g\in X_{n}}\|f-g\|,\quad f\in K.$ (3.2)
The constant $\gamma_{n}:=\|I-\Pi_{n}\|_{K^{\prime}}$ is usually bounded as
$\gamma_{n}\leq 1+\|\Pi_{n}\|_{K^{\prime}}=1+\Lambda_{n}.$ (3.3)
The formula (2.2) implies the upper bound $\Lambda_{n}\leq\big{(}\max_{1\leq i\leq n}\|\ell_{i}\|_{K^{\prime}}\big{)}\sum_{j=1}^{n}\|h_{j}\|$. Then the iterate $f_{n}$ in Step 2 of Algorithm 2.2 satisfies
$\inf_{g\in X_{n-1}}\|f_{n}-g\|={\rm dist}(f_{n},X_{n-1})\geq\gamma_{n-1}^{-1}\sup_{f\in K}{\rm dist}(f,X_{n-1}),$ (3.4)
the same as the selection criterion in the weak greedy RBM (Algorithm 2.1).
Let $\ell_{2}^{n}$ denote the Euclidean space $\mathbb{R}^{n}$ under the $\ell_{2}$-norm. Our analysis utilizes the following quantity (see Section III.B. of [39])
$\delta_{n}=\delta_{X,n}:=\sup_{\begin{subarray}{c}Y_{n}\subset X\\ \dim Y_{n}=n\end{subarray}}\inf_{T:Y_{n}\rightarrow\ell_{2}^{n}}\|T\|_{Y_{n}\rightarrow\ell_{2}^{n}}\|T^{-1}\|_{\ell_{2}^{n}\rightarrow Y_{n}},$
where the infimum is taken over all isomorphism $T$ from the Banach space $Y_{n}$ to $\ell_{2}^{n}$. In fact, $\inf_{T:Y_{n}\rightarrow\ell_{2}^{n}}\|T\|_{Y_{n}\rightarrow\ell_{2}^{n}}\|T^{-1}\|_{\ell_{2}^{n}\rightarrow Y_{n}}$ is called the Banach–Mazur distance between $Y_{n}$ and $\ell_{2}^{n}$. Our analysis of the EIM and the CGA relies on the next fundamental lemma, which was developed in [18].
Lemma 3.1 (Lemma 3.1 from [18]).
Let $K$ be a compact set in a Banach space $X$. Let $V_{n}=\pi^{\frac{n}{2}}/\Gamma(\frac{n}{2}+1)$ be the volume of an $\ell_{2}$ unit ball in $\mathbb{R}^{n}$. For any $v_{1},\ldots,v_{n}\in K$ with $X_{k}={\rm span}\{v_{1},\ldots,v_{k}\}$ and $X_{0}=\{0\}$, we have
$\Big{(}\prod_{k=1}^{n}{\rm dist}(v_{k},X_{k-1})\Big{)}^{\frac{1}{n}}\leq\delta_{X,n}(n!V_{n})^{\frac{1}{n}}\varepsilon_{n}({\rm co}(K)).$
Now we are in a position to present our convergence estimates of the generalized EIM.
Theorem 3.2.
For the weak greedy RBM (Algorithm 2.1), we have
$\sup_{f\in K}{\rm dist}(f,X_{n-1})\leq(\alpha_{1}\cdots\alpha_{n})^{-\frac{1}{n}}{\delta}_{n}(n!V_{n})^{\frac{1}{n}}\varepsilon_{n}({\rm co}(K)).$
In particular, the generalized EIM (Algorithm 2.2) satisfies
$\sup_{f\in K}\|f-\Pi_{n-1}f\|\leq\gamma_{n-1}(\gamma_{1}\cdots\gamma_{n-1})^{\frac{1}{n}}{\delta}_{n}(n!V_{n})^{\frac{1}{n}}\varepsilon_{n}({\rm co}(K)).$
Proof 3.3.
First we focus on Algorithm 2.1. It follows from Step 1 of Algorithm 2.1 and Lemma 3.1 with $v_{k}=f_{k}$ in Algorithm 2.1 that
$\sup_{f\in K}{\rm dist}(f,X_{n-1})\leq\Big{(}\prod_{k=1}^{n}\sup_{f\in K}{\rm dist}(f,X_{k-1})\Big{)}^{\frac{1}{n}}$
$\leq\Big{(}\prod_{k=1}^{n}\alpha_{k}^{-1}{\rm dist}(f_{k},X_{k-1})\Big{)}^{\frac{1}{n}}$
$\leq(\alpha_{1}\cdots\alpha_{n})^{-\frac{1}{n}}\delta_{n}(n!V_{n})^{\frac{1}{n}}\varepsilon_{n}({\rm co}(K)).$
Next we note that $\{f_{n}\}_{n\geq 1}$ in Algorithm 2.2 satisfies (3.4), the same selection criterion in the weak greedy RBM (Algorithm 2.1) with $\alpha_{n}=\gamma_{n-1}^{-1}$ and $\gamma_{0}=1$. Combining (3.2) and the first part of this theorem completes the proof.
It is easy to see that $\delta_{X,n}=1$ provided $X$ is a Hilbert space. For Banach spaces and Sobolev spaces $W^{k}_{p}(\Omega)$, it has been proven in Section III.B.9 of [39] that
${\delta}_{X,n}\leq\left\{\begin{aligned} &\sqrt{n},\quad\text{ when $X$ is a general Banach space},\\ &n^{|\frac{1}{2}-\frac{1}{p}|},\quad\text{when $W^{k}_{p}(\Omega)$ with $p\in[1,\infty]$},\end{aligned}\right.$ (3.5)
where the convention $L_{p}(\Omega)=W^{0}_{p}(\Omega)$ is adopted. We recall that $C$ is an absolute constant independent of $n$. With the help of (3.5), several important asymptotic consequences of Theorem 3.2 are summarized in the next corollary.
Corollary 3.4.
When $X$ is a general Banach space, we have
$\sup_{f\in K}\|f-\Pi_{n-1}f\|\leq C(1+\Lambda_{n-1})\left(\prod_{k=1}^{n-1}(1+\Lambda_{k})\right)^{\frac{1}{n}}n\varepsilon_{n}({\rm co}(K)).$ (3.6)
When $X=W^{k}_{p}(\Omega)$ with $p\in[1,\infty]$, it holds that
$\sup_{f\in K}\|f-\Pi_{n-1}f\|\leq C(1+\Lambda_{n-1})\left(\prod_{k=1}^{n-1}(1+\Lambda_{k})\right)^{\frac{1}{n}}n^{\frac{1}{2}+|\frac{1}{2}-\frac{1}{p}|}\varepsilon_{n}({\rm co}(K)).$ (3.7)
When $X$ is a Hilbert space, it holds that
$\sup_{f\in K}\|f-\Pi_{n-1}f\|\leq C\Lambda_{n-1}\left(\Lambda_{1}\cdots\Lambda_{n-1}\right)^{\frac{1}{n}}n^{\frac{1}{2}}\varepsilon_{n}({\rm co}(K)).$ (3.8)
Proof 3.5.
The estimates (3.6) and (3.7) follow from Theorem 3.2, the simple bound $\gamma_{n}\leq 1+\Lambda_{n}$, (3.5) and the well-known formulae
[…]
The third estimate (3.8) is a result of Theorem 3.2 and the classical identity
$\gamma_{n}=\|I-\Pi_{n}\|_{K^{\prime}}=\|\Pi_{n}\|_{K^{\prime}}=\Lambda_{n}$
(see [41]) for the idempotent ($\Pi_{n}^{2}=\Pi_{n}$) operator $\Pi_{n}$ in an inner-product space.
3.1 Convergent Examples of the EIM
Recall that $\mathcal{P}$ is a compact set in $\mathbb{R}^{p}$. In practice, the EIM is applied to the parametrized function set
$K=\{g_{\mu}\in X:\mu\in\mathcal{P}\subset\mathbb{R}^{d}\}.$
When the mapping $\mu\mapsto g_{\mu}$ is of smoothness order $\alpha>0$ (see [33] for the accurate definition), it is known that
$d_{n}(K)\leq Cn^{-\frac{\alpha}{d}}.$ (3.9)
In addition, if $X$ is a type-2 Banach space, e.g., $X=W^{k}_{p}(\Omega)$ with $2\leq p<\infty$, it has been shown in [33] that
$\varepsilon_{n}({\rm co}(K))\leq Cn^{-\frac{1}{2}-\frac{\alpha}{d}}.$ (3.10)
Under the condition (3.9), the classical $n$-width based convergence result (1.1) from [24] yields
$\sup_{f\in K}\|f-\Pi_{n}f\|\leq C(1+\Lambda_{n})^{3}n^{-\frac{\alpha}{d}+1}.$ (3.11)
On the other hand, it follows from (3.10) and Corollary 3.4 with $X=W^{k}_{p}(\Omega)$ and $1<p<\infty$ that
$\sup_{f\in K}\|f-\Pi_{n-1}f\|\leq C(1+\Lambda_{n-1})\left(\prod_{k=1}^{n-1}(1+\Lambda_{k})\right)^{\frac{1}{n}}n^{-\frac{\alpha}{d}+|\frac{1}{2}-\frac{1}{p}|},$ (3.12)
improving the existing result (3.11). Compared with classical error analysis, our error estimate (3.12) depends only mildly on the Lebesgue constant $\Lambda_{n}$. For common applications, numerical evidence (see [25]) shows that $\Lambda_{n}=O(n^{\beta})$ with $\beta>0$ and thus confirms another advantage of (3.12) over classical error bound (3.11).
In the end of this section, we remark that the upper bound (3.3) for $\gamma_{n}$ turns out to be improvable in many Banach spaces. In fact, the work [34] generalized the identity in [41] from Hilbert spaces to Banach spaces, see the next lemma.
Lemma 3.6 (Theorem 3 from [34]).
In a normed space $Y$, let $P:Y\rightarrow Z\subseteq Y$ be a linear projection ($P^{2}=P$) onto a finite-dimensional subspace $Z$. Then we have
$\|I-P\|_{Y\rightarrow Z}\leq\min\big{(}1+\|P\|_{Y\rightarrow Z}^{-1},\delta_{Y,2}^{2}\big{)}\|P\|_{Y\rightarrow Z}.$
Let $C_{n}:=\min\big{(}1+\Lambda_{n}^{-1},\delta_{X,2}^{2}\big{)}$. Lemma 3.6 with $P=\Pi_{n}$ implies that
$\gamma_{n}=\|I-\Pi_{n}\|_{K^{\prime}}\leq C_{n}\|\Pi_{n}\|_{K^{\prime}}=C_{n}\Lambda_{n},$ (3.13)
and thus slightly improve the bound (3.3) when $\delta_{X,2}$ is small enough. Due to Theorem 3.2 and (3.13), in the case $X=W^{k}_{p}(\Omega)$ ($\delta_{X,2}\leq 2^{|\frac{1}{2}-\frac{1}{p}|}$), Corollary 3.4 is slightly improved as
$\sup_{f\in K}\|f-\Pi_{n-1}f\|\leq CC_{n-1}\big{(}C_{1}\cdots C_{n-1}\big{)}^{\frac{1}{n}}n^{\frac{1}{2}+|\frac{1}{2}-\frac{1}{p}|}\varepsilon_{n}({\rm co}(K)),$
where $C_{n}=\min\{1+\Lambda_{n}^{-1},2^{|1-\frac{2}{p}|}\}$.
4 Convergence of the CGA
We are going to derive convergence estimates of the CGA (Algorithm 2.4) in terms of the metric entropy of ${\rm co}(K)$. Another key concept used in our analysis is the modulus of smoothness of the Banach space $X$:
$\rho_{X}(t)=\sup_{\|x\|=1,\|y\|=t}\left\{\frac{1}{2}\|x+y\|+\frac{1}{2}\|x-y\|-1\right\}.$ (4.1)
It is straightforward to check that $\rho_{X}(t)\leq t$. In this section, we assume that for some $s\in(1,2]$ and $t>0$, $\rho_{X}$ satisfies
$\rho_{X}(t)\leq C_{X}t^{s},$ (4.2)
where $C_{X}>0$ is a constant. In fact, the convergence rate of $\rho_{X}$ as $t\to 0^{+}$ is known for many Banach spaces. For example, for $t>0$ we have (see Section 1.e. of [20])
$\rho_{X}(t)\leq\left\{\begin{aligned} &C_{X}t^{p},\quad\text{when }X=W^{k}_{p}(\Omega),~{}1\leq p<2,\\ &C_{X}t^{2},\quad\text{when }X=W^{k}_{p}(\Omega),~{}2\leq p<\infty.\end{aligned}\right.$ (4.3)
Under our assumption (4.2), the peak functional $F_{r_{n}}$ in Algorithm 2.4 satisfies
$F_{r_{n}}(\varphi)=0,\quad\varphi\in X_{n}.$ (4.4)
In fact, (4.4) also holds when $X$ is only a uniformly smooth Banach space ($\rho_{X}(t)=o(t)$ as $t\to 0^{+}$), see Lemma 2.1 in [36]. For the target function $f\in X$, we define
$\|f\|_{\mathcal{L}_{1}(K)}:=\inf\Big{\{}\sum_{i}|c_{i}|:f=\sum_{i}c_{i}g_{i},~{}g_{i}\in K\text{ for each }i\Big{\}}.$
For example, we have $\|f\|_{\mathcal{L}_{1}(K)}\leq c$ when $f/c\in{\rm co}(K)$. The next theorem is our main result about the convergence of the weak CGA.
Theorem 4.1.
Assume (4.2) holds. For the weak CGA (Algorithm 2.4) we have
$\|f-f_{n}\|\leq 2^{1+\frac{1}{s}}C_{X}^{\frac{1}{s}}(\alpha_{1}\cdots\alpha_{n})^{-\frac{1}{n}}\delta_{n}n^{\frac{1}{s}-1}(n!V_{n})^{\frac{1}{n}}\|f\|_{\mathcal{L}_{1}(K)}\varepsilon_{n}({\rm co}(K)).$
Proof 4.2.
Let $\bar{g}_{n}\in X_{n-1}$ be the element in $X_{n-1}$ that is closest to $g_{n}$ in Algorithm 2.4. For convenience, we set
$\phi_{n}=g_{n}-\bar{g}_{n},\quad\|\phi_{n}\|={\rm dist}(g_{n},X_{n-1}).$ (4.5)
The definition of $\rho_{X}$ (4.1) implies for any $x,y\in X,$
$\|x+y\|+\|x-y\|\leq 2\|x\|\left(\rho\left(\frac{\|y\|}{\|x\|}\right)+1\right).$
[…]
$\|r_{n-1}-s_{n}\lambda\phi_{n}\|+\|r_{n-1}+s_{n}\lambda\phi_{n}\|$ (4.6)
$\leq 2\|r_{n-1}\|\left(1+\rho\left(\frac{\lambda\|\phi_{n}\|}{\|r_{n-1}\|}\right)\right).$
Using the definition of $F_{r_{n-1}}$ and $g_{n}$ in Algorithm 2.4 and (4.4), (2.3), we have
$\alpha_{n}^{-1}|F_{r_{n-1}}(\phi_{n})|=\alpha_{n}^{-1}|F_{r_{n-1}}(g_{n})|\geq\sup_{g\in K}|F_{r_{n-1}}(g)|$ (4.7)
$\geq\|f\|^{-1}_{\mathcal{L}_{1}(K)}F_{r_{n-1}}(f)=\|f\|^{-1}_{\mathcal{L}_{1}(K)}F_{r_{n-1}}(r_{n-1})=\|f\|^{-1}_{\mathcal{L}_{1}(K)}\|r_{n-1}\|.$
As a result of (4.7),
$\|r_{n-1}+\lambda s_{n}\phi_{n}\|\geq F_{r_{n-1}}(r_{n-1}+\lambda s_{n}\phi_{n})$ (4.8)
$=\|r_{n-1}\|+\lambda|F_{r_{n-1}}(\phi_{n})|\geq\Big{(}1+\lambda\alpha_{n}\|f\|^{-1}_{\mathcal{L}_{1}(K)}\Big{)}\|r_{n-1}\|.$
Combining (4.6) with (4.8) then leads to
$\|r_{n}\|\leq\inf_{\lambda>0}\|r_{n-1}-\lambda s_{n}\phi_{n}\|$ (4.9)
$\leq 2\|r_{n-1}\|\left(1+\rho\left(\frac{\lambda\|\phi_{n}\|}{\|r_{n-1}\|}\right)\right)-\Big{(}1+\lambda\alpha_{n}\|f\|^{-1}_{\mathcal{L}_{1}(K)}\Big{)}\|r_{n-1}\|$
$\leq\|r_{n-1}\|\left(1-\lambda\alpha_{n}\|f\|^{-1}_{\mathcal{L}_{1}(K)}+2C_{X}\left(\frac{\lambda\|\phi_{n}\|}{\|r_{n-1}\|}\right)^{s}\right).$
Now we set $t=s/(s-1)>1$ and
$2C_{X}\left(\frac{\lambda\|\phi_{n}\|}{\|r_{n-1}\|}\right)^{s}=\frac{\lambda}{2}\alpha_{n}\|f\|_{\mathcal{L}_{1}(K)}^{-1}.$
Direct calculation shows that
$\lambda=\|\phi_{n}\|^{-t}\|r_{n-1}\|^{t}\alpha_{n}^{\frac{1}{s-1}}\Big{(}4C_{X}\|f\|_{\mathcal{L}_{1}(K)}\Big{)}^{-\frac{1}{s-1}}.$
Then (4.9) reduces to
$\|r_{n}\|\leq\|r_{n-1}\|\big{(}1-2^{-1}(4C_{X})^{-\frac{1}{s-1}}\alpha_{n}^{t}\|\phi_{n}\|^{-t}\|f\|_{\mathcal{L}_{1}(K)}^{-t}\|r_{n-1}\|^{t}\big{)}.$ (4.10)
We define $a_{n}=\|r_{n}\|/\|f\|_{\mathcal{L}_{1}(K)}$, $b_{n}=2^{-1}(4C_{X})^{-\frac{1}{s-1}}\alpha_{n}^{t}\|\phi_{n}\|^{-t}$. Taking the $t$-th power on both sides of (4.10) leads to the recurrence relation
$a_{n}^{t}\leq a_{n}^{t}\big{(}1-b_{n}a_{n}^{t}\big{)}.$ (4.11)
Using (4.11), the induction lemma (Lemma 4.1) in [18] and Lemma 3.1, we obtain
$a_{n}^{t}$ $\leq\frac{1}{1+b_{1}+\cdots+b_{n}}\leq\frac{1}{n}\left(b_{1}\cdots b_{n}\right)^{-\frac{1}{n}}$
$=2(4C_{X})^{\frac{1}{s-1}}(\alpha_{1}\cdots\alpha_{n})^{-\frac{t}{n}}\frac{1}{n}\left(\prod_{k=1}^{n}{\rm dist}(g_{k},X_{k-1})\right)^{\frac{t}{n}}$
$\leq 2(4C_{X})^{\frac{1}{s-1}}(\alpha_{1}\cdots\alpha_{n})^{-\frac{t}{n}}\delta_{n}^{t}\frac{(n!V_{n})^{\frac{t}{n}}}{n}\varepsilon_{n}({\rm co}(K))^{t}.$
The proof is complete.
Convergence analysis of the weak CGA can be extended to a target function $f$ without bounded $\|\bullet\|_{\mathcal{L}_{1}(K)}$ norm. The key tool is the interpolation space based on the $K$-functional
$K(t,g):=\inf_{h\in\mathcal{L}_{1}(K)}\left(\|g-h\|_{X}+t\|h\|_{\mathcal{L}_{1}(K)}\right),$
where $t>0$ and $g\in X$ (see [9]). For the index $\theta\in(0,1)$, the interpolation norm between $X$ and $\mathcal{L}_{1}(K)$ is
$\|g\|_{\theta}=\|g\|_{[X,\mathcal{L}_{1}(K)]_{\theta,\infty}}:=\sup_{0<t<\infty}t^{-\theta}K(t,g).$
Then the interpolation space
$X_{\theta}:=\{f\in X:\|f\|_{\theta}<\infty\}$
is larger than $\mathcal{L}_{1}(K)$. Following the same analysis in [2], we obtain an explicit and improved error estimate of the weak CGA for $f\in X_{\theta}$.
Corollary 4.3.
For all $f\in X_{\theta}$ and $\theta\in(0,1)$, Algorithm 2.4 satisfies
$\|f-f_{n}\|\leq 2^{\theta+\frac{\theta}{s}}C_{X}^{\frac{\theta}{s}}(\alpha_{1}\cdots\alpha_{n})^{-\frac{\theta}{n}}\delta^{\theta}_{n}n^{\frac{\theta}{s}-\theta}(n!V_{n})^{\frac{\theta}{n}}\|f\|_{\theta}\varepsilon_{n}({\rm co}(K))^{\theta}.$
4.1 Convergent Examples of the CGA
We recall the classical convergence result of the weak CGA given in [36]:
$\|f-f_{n}\|\leq C(1+\alpha_{1}^{\frac{s}{s-1}}+\cdots+\alpha^{\frac{s}{s-1}}_{n})^{\frac{1}{s}-1},\quad f\in{\rm co}(K).$ (4.12)
Combining (4.12) with (4.3) then yields
$\|f-f_{n}\|\leq\left\{\begin{aligned} &C(1+\alpha_{1}^{2}+\cdots+\alpha^{2}_{n})^{\frac{1}{2}}\quad\text{ when }X=W^{k}_{p}(\Omega),~{}2\leq p<\infty,\\ &C(1+\alpha_{1}^{\frac{p}{p-1}}+\cdots+\alpha^{\frac{p}{p-1}}_{n})^{\frac{1}{p}-1}\quad\text{ when }X=W^{k}_{p}(\Omega),~{}1<p<2.\end{aligned}\right.$ (4.13)
In contrast, Theorem 4.1 with $(n!V_{n})^{\frac{1}{n}}=O(n^{\frac{1}{2}})$ yields the asymptotic estimates
$\|f-f_{n}\|\leq C(\alpha_{1}\cdots\alpha_{n})^{-\frac{1}{n}}\delta_{n}n^{\frac{1}{s}-\frac{1}{2}}\|f\|_{\mathcal{L}_{1}(K)}\varepsilon_{n}({\rm co}(K)).$ (4.14)
We note that the threshold constants $\{\alpha_{i}\}_{i\geq 1}$ enter our error bound in a multiplicative way. It follows from (4.14) and (4.3), (3.5) that
$\|f-f_{n}\|\leq\left\{\begin{aligned} &C(\alpha_{1}\cdots\alpha_{n})^{-\frac{1}{n}}n^{\frac{1}{2}-\frac{1}{p}}\|f\|_{\mathcal{L}_{1}(K)}\varepsilon_{n}({\rm co}(K))\\ &\qquad\text{ when }X=W^{k}_{p}(\Omega),~{}2\leq p<\infty,\\ &C(\alpha_{1}\cdots\alpha_{n})^{-\frac{1}{n}}n^{\frac{2}{p}-1}\|f\|_{\mathcal{L}_{1}(K)}\varepsilon_{n}({\rm co}(K))\\ &\qquad\text{ when }X=W^{k}_{p}(\Omega),~{}1<p<2.\end{aligned}\right.$ (4.15)
Theorem 4.1 and its consequences (4.14), (4.15) are favorable when $\varepsilon_{n}({\rm co}(K))$ is small enough.
To illustrate the advantage of Theorem 4.1 over (4.12), we consider the $\texttt{ReLU}_{m}$ activation function $\sigma_{m}$ defined by $\sigma_{m}(t)=\max(t^{m},0)$ with $m\geq 1$ and $\sigma_{0}(t)=\max\{{\rm sign}(t),0\}$ for $t\in\mathbb{R}$. On $\Omega\subset\mathbb{R}^{d}$, the following dictionary
$K=\big{\{}\sigma_{m}(\omega\cdot\bullet+b)\in L_{p}(\Omega):\omega\in\mathbb{R}^{d},~{}\|\omega\|_{\ell^{2}}=1,~{}\beta_{0}\leq b\leq\beta_{1}\big{\}}$ (4.16)
is widely used in machine learning by shallow neural networks, where $\beta_{0}$, $\beta_{1}$ are constants depending on $\Omega$. When $\Omega=B_{1}^{d}$ is a $d$-dimensional unit ball, it is proved in Theorem 3 of [31] and Theorem 10 of [33] that
$\varepsilon_{n\log n}({\rm co}(K))_{L_{p}(B_{1}^{d})}\leq Cn^{-\frac{1}{2}-\frac{2m+1}{2d}}.$ (4.17)
Therefore, for the CGA with $\alpha_{1}=\cdots=\alpha_{n}=1$ and $X=L_{p}(\Omega)$, the proposed error bound (4.15) improves upon the existing result (4.13) provided $\frac{2m+1}{2d}>|\frac{1}{2}-\frac{1}{p}|$.
4.2 Numerical Convergence of the CGA
In the Banach space $L_{p}(\Omega)$ on $\Omega=[0,1]$, we test the convergence of the CGA based on the dictionary (4.16) with $m=0$, $\beta_{0}=-2$, $\beta_{1}=2$. The target function is set to be $f(x)=\sin(\pi x)$ for $x\in\Omega$. For any $g\in L_{p}(\Omega)$, its peak functional $F_{g}\in L_{\frac{p}{p-1}}(\Omega)$ is
$F_{g}=\text{sign}(g)|g|^{p-1}/\|g\|^{p-1}_{L_{p}(\Omega)}.$
The convergence history of the CGA is displayed in Figure 1. We note the estimate (4.15) with (4.17) ensures that
$\|f-f_{\lceil{n\log n}\rceil}\|_{L_{p}(\Omega)}\leq Cn^{-\frac{1}{p}-\frac{1}{2}}\|f\|_{\mathcal{L}_{1}(K)},\quad p\geq 2.$
However, it is observed from Figure 1 that the convergence rate of the CGA in $L_{p}$ is approximately $O(n^{-1})$ regardless of the value of $p$.
Currently we are not sure whether the convergence rate in Theorem 4.1 is sharp or not. Such questions seem to be technical and open even for the classical convergence result (4.12). Therefore, verifying the sharpness of the proposed convergence rate of CGAs would be a future research direction.
Figure 1: Convergence history of the CGA error, $\log(E)=\log_{10}\|r_{n}\|_{L_{p}(\Omega)}$.